A wave is travelling at a speed of 360 meterper second with frequency 400 hz . The points on the waves are 0.2 meter apart find the phase difference between them in the direction of travelling?

Grade:12th Pass

1 Answers

Arun
6 years ago
Dear Abdul
 
Lambda = 360/400 = 0.9
 
Now
Phase difference (phi) = (0.2/0.9 ) * 2*pi = 4*pi/9
 
 
Regards
Arun (askIITians forum expert)